OpenStudy (anonymous):

derivative of sqrt(2x + 1)

OpenStudy (danjs):

= derivative of (2x+1)^(1/2)

OpenStudy (anonymous):

Dan, can you explain why the first line of your answer is such?

OpenStudy (danjs):

Substitute u = 2x+1

OpenStudy (danjs):

\[\frac{ dy }{ dx }=\frac{ dy }{ du }*\frac{ du }{ dx }\]

OpenStudy (danjs):

Substitute u = 2x+1 so you have derivative of sqrt(u)

OpenStudy (danjs):

then follow the chain rule as i stated above. find dy/du and multiply it by du/dx u = 2x+1 du=2 dx du/dx=2

OpenStudy (danjs):

Derivative of sqrt u = (1/2)u^(-1/2)

OpenStudy (danjs):

so \[\frac{ 1 }{ 2 }\frac{ 1 }{ \sqrt{u} }*\frac{ du }{ dx } = \frac{ 1 }{ 2 }\frac{ 1 }{ \sqrt{u} }*2\]

OpenStudy (danjs):

now resubstitute back in for u=2x+1

OpenStudy (danjs):

1/sqrt u

OpenStudy (anonymous):

thanks for the great explaination, watching some videos on Kahn to better-understand the process

OpenStudy (danjs):

just remember for a parenthesis to a power just take the derivative of the quantity as a whole, then multiply that by the derivative of the inside of the parenthesis after doing a couple you will get it down pretty easy you dont have to do the whole substitution thing in your head